OpenStudy (anonymous):

can someone derive this for me? f(x)=sin (cos (5x))

jimthompson5910 (jim_thompson5910):

f(x)=sin (cos (5x)) f'(x)=cos (cos (5x))*d/dx[cos(5x)] f'(x)=cos (cos (5x))*(-sin(5x))*d/dx[5x] f'(x)=cos (cos (5x))*(-sin(5x))*5x f'(x)=-5x*cos (cos (5x)*sin(5x)

OpenStudy (anonymous):

ok, got it! thanks!! I had something similar, but missed a step...